# Catalyst4D: High-Fidelity 3D-to-4D Scene Editing via Dynamic Propagation

### 13 Mar 2026 v1

@@ 2 Related Work
The introduction of 3DGS [18, 26, 27, 28] has addressed many of these limitations by enabling point-based scene representations with efficient rendering and flexible editing. Methods such as GaussiansEditors [7, 47] constrain gradient-based optimization to target 3D Gaussians, achieving efficient object-level editing while leveraging 2D diffusion models as priors. Subsequent works [6, 52, 48, 49] incorporate 3D structural information into diffusion-based editing to improve cross-view consistency, and others [1, 53, 10, 56] exploit depth information from 3DGS for geometry modification. High-quality stylization methods [29, 11, 59] extend image style transfer techniques to 3D scenes, further broadening 3DGS applications in artistic creation. Despite their capabilities, these approaches remain confined to static scenes. Building upon these strong 3D editing foundations, extending them to dynamic scenes with temporal coherence remains an open and compelling challenge.
@@ 3 Scene Representation
3D Gaussian Splatting. We adopt anisotropic 3D Gaussians [18] to represent the scene. Each Gaussian is defined by a center $\bm{\mu}\in\mathbb{R}^{3}$ and a covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$, whose spatial influence at a point $\mathbf{x}$ is given by:
$G(\mathbf{x})=\exp\!\left(-\tfrac{1}{2}(\mathbf{x}-\bm{\mu})^{\mathrm{T}}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\right).$ (1)
The covariance is factorized as $\bm{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\mathrm{T}}\mathbf{R}^{\mathrm{T}}$, where $\mathbf{R}$ is a rotation matrix (parameterized by quaternion $\mathbf{q}$) and $\mathbf{S}$ is a scaling matrix (from scale vector $\mathbf{s}$). Each Gaussian additionally has an opacity $\sigma$ and spherical harmonics coefficients $\mathbf{sh}$ to encode view-dependent color. The full scene can thus be represented as a set: $\mathcal{G}=\{G_{i}:\bm{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i},\sigma_{i},\mathbf{sh}_{i}\}_{i=1}^{N}$. During differentiable rendering, Gaussians are projected to 2D and composited via $\alpha$-blending. The pixel color $C$ along ray $\mathbf{o}$ is:
$C(\mathbf{o})=\sum_{i\in\mathcal{N}}\operatorname{SH}(\mathbf{sh}_{i},\mathbf{v})\alpha_{i}~\!\!\!\prod_{j=1}^{i-1}\!\!~(1-\alpha_{j}),$ (2)
@@
$\Delta\bm{\mu},\Delta\mathbf{q},\Delta\mathbf{s}=\mathcal{F}_{\theta}(\bm{\mu}_{c},t;\mathbf{z}),$ (3)
where $\mathbf{z}$ is a scene-level latent code. The deformed parameters at time $t$ are computed as: $\bm{\mu}_{t}=\bm{\mu}_{c}+\Delta\bm{\mu}$, $\mathbf{q}_{t}=\mathbf{q}_{c}+\Delta\mathbf{q}$, and $\mathbf{s}_{t}=\mathbf{s}_{c}+\Delta\mathbf{s}$. This deformation-based formulation ensures temporal coherence and semantic consistency across frames. Our framework is compatible with various 4D Gaussian formulations, and in practice, we build upon Swift4D [51] and 4DGS [50], which serve as the base representations in our experiments.
@@ 4 Method
As illustrated in Fig. 2, Catalyst4D transfers 3D Gaussian edits from the canonical frame to the entire 4D sequence. Given an edited first-frame Gaussians obtained by existing 3D Gaussian editors (e.g.[6, 19, 11]), our goal is to propagate the edits through time while maintaining geometric accuracy and temporal coherence. To this end, Catalyst4D introduces two key components: (1) Anchor-based Motion Guidance (AMG), which establishes region-level correspondences via anchors to enable localized, consistent deformation transfer; and (2) Color Uncertainty-guided Appearance Refinement (CUAR), which corrects occlusion-induced color artifacts by propagating confident appearance information along motion trajectories.
@@ 4.1 Anchor-based Motion Guidance
To establish correspondences between the original and edited anchors, we compute a soft correspondence matrix $P\in\mathbb{R}^{n\times m}$ using unbalanced optimal transport (UOT) [3, 43], solved via the Sinkhorn algorithm [8]. Here, each element $P_{ij}$ represents the transport from edited anchor $\mathbf{p}_{j}^{\mathrm{edit}}$ to source anchor $\mathbf{p}_{i}^{\mathrm{src}}$, and reliable correspondences are written as:
$corr=\{(\mathbf{p}_{i}^{\mathrm{src}},\mathbf{p}_{j}^{\mathrm{edit}})\mid i=\operatorname*{arg\,max}_{k}P_{kj}\}.$ (8)
@@
Specifically, at timestamp $t$, the position deformation $\Delta\bm{\mu}_{\mathbf{g}}^{t}$ is computed as a weighted average of the corresponding source Gaussians’ deformations:
$\Delta\bm{\mu}_{\mathbf{g}}^{t}=\frac{\sum_{\mathbf{g}^{\prime}\in\mathcal{G}^{\mathrm{sub}}_{\mathrm{src}}}w_{\mathbf{g}^{\prime}}\Delta\bm{\mu}_{\mathbf{g}^{\prime}}^{t}}{\sum_{\mathbf{g}^{\prime}\in\mathcal{G}^{\mathrm{sub}}_{\mathrm{src}}}w_{\mathbf{g}^{\prime}}},$ (9)
where $\Delta\bm{\mu}_{\mathbf{g}^{\prime}}^{t}$ is the temporal deformation of source Gaussian $\mathbf{g}^{\prime}$ from frame 1 to $t$, predicted by the deformation network. The weight $w_{\mathbf{g}^{\prime}}$ combines the Gaussian’s opacity $\sigma_{\mathbf{g}^{\prime}}$ and its Mahalanobis distance to $\mathbf{g}$:
$w_{\mathbf{g}^{\prime}}=\sigma_{\mathbf{g}^{\prime}}\exp\Big(-\frac{1}{2}(\bm{\mu}_{\mathbf{g}^{\prime}}-\bm{\mu}_{\mathbf{g}})^{\mathrm{T}}\bm{\Sigma}_{\mathbf{g}^{\prime}}^{-1}(\bm{\mu}_{\mathbf{g}^{\prime}}-\bm{\mu}_{\mathbf{g}})\Big).$ (10)
@@ 4.2 Color Uncertainty-guided Appearance Refinement
Optical Flow Rendering. Given the temporal deformation $\Delta\bm{\mu}^{t}$ from motion guidance, we first render optical flow maps describing the image-space displacement of each Gaussian from frame 1 to frame $t$. Inspired by [20, 61], for a viewpoint $v$, the 3D positions $\bm{\mu}_{1}$ and $\bm{\mu}_{t}$ are projected to 2D via the projection function $\operatorname{Proj}_{v}$, yielding the per-Gaussian displacement:
$f_{1\to t}^{v}=\operatorname{Proj}_{v}(\bm{\mu}_{t})-\operatorname{Proj}_{v}(\bm{\mu}_{1}).$ (11)
@@ 5.1 Experimental Setup
Dataset and Metrics. We conduct experiments on the multi-camera video datasets DyNeRF [25] and MeetRoom [23], as well as the monocular dataset HyperNeRF [35]. We compare our method with state-of-the-art approaches: Instruct 4D-to-4D [34], CTRL-D [13], and Instruct-4DGS [22]. The editing results of all methods are evaluated on test views using CLIP [37] similarity and a temporal consistency score [15].
Implementation Details. We incorporate several baseline 3D editing methods, including the 3D-consistent editing approach DGE [6], the SDS inversion-based framework DreamCatalyst [19], and the high-resolution style transfer method SGSST [11]. For 4D scene representation, We employ Swift4D [51] under multi-camera settings, while utilizing 4DGS [50] for monocular scenes. Following the settings of 4DGS [50], we constrain the edited dynamic Gaussians to share consistent opacity and color attributes across all frames during optimization.
@@ 5.2 Qualitative Results
Qualitative comparison. We conduct text-guided editing across all methods to ensure fair comparison. The qualitative results achieved on different datasets are shown in Fig. 4, with the corresponding text prompts displayed on the leftmost side. As can be seen, Instruct 4D-to-4D [34] and Instruct-4DGS [22] lacks fine-grained localization capabilities, often causing unintended modifications to non-target regions. Although CTRL-D [13] supports localized editing in dynamic scenes, its reliance on 2D image edits introduces limitations. When the 2D edits deviate significantly from the original scene content, the reconstructed 4D representation fails to accurately reflect them, resulting in degraded visual fidelity (see the 3rd and 5th rows of CTRL-D in Fig.4). In contrast, our 3D editing component employs iterative alignment of 2D and 3D representations, producing more natural and coherent edits. Building on these geometry-consistent 3D edits, Catalyst4D focuses on propagating them to 4D Gaussian scenes, achieving enhanced spatio-temporal consistency and visual realism. Furthermore, although the compared CTRL-D appears visually plausible, it still introduces undesired modifications in non-edited regions due to the limitations of the underlying 2D diffusion model. As illustrated in Fig.5, while CTRL-D edits the character, it also causes noticeable deviations in non-target objects such as the dog on the stool and objects on the table compared to the original scene. By contrast, our method constrains target dynamic Gaussians directly through 3D editing gradients, enabling more precise and localized editing in dynamic scenes.
Figure 5: Qualitative comparison of localized editing. In contrast to CTRL-D, which introduces inconsistencies in non-edited regions, our method achieves more precise and localized editing by constraining dynamic Gaussians via 3D editing gradients.
@@ 5.3 Quantitative Results
Tab. 1 presents quantitative comparisons with Instruct 4D-to-4D (IN4D), Instruct-4DGS (I4DGS), and CTRL-D on two DyNeRF scenes (Sear-steak, Coffee-martini) and one MeetRoom scene (Trimming). Across all three scenes, our method achieves the highest CLIP similarity, indicating the strongest semantic alignment with the editing instructions. Regarding efficiency, our method outperforms CTRL-D and Instruct 4D-to-4D. In terms of temporal coherence, measured by Consistency from VBench [15], our method also delivers consistently strong performance, matching or surpassing CTRL-D in most cases. We observe a slight drop in consistency on only one scene, which stems from our design choice of not retraining the deformation network or modifying the density of edited dynamic Gaussians. Despite this, the overall results demonstrate that our method maintains competitive stability while offering significantly better semantic fidelity.
5.4 Ablation Studies
In contrast, AMG module establishes robust structural correspondences between edited and original Gaussians by matching sparse yet representative anchor points. This preserves regional motion and prevents interference from sparse noise, leading to improved geometric fidelity. The quantitative results in Tab. 2 confirm this observation, our method achieves higher CLIP similarity and better temporal consistency than relying on the deformation network to propagate the edited Gaussians.
@@ 6 Conclusion
We present Catalyst4D, a novel framework that brings the editing capabilities of 3D methods into the realm of dynamic 4D Gaussian scenes. By introducing an anchor-based motion guidance strategy to preserve structural consistency and a color uncertainty-guided appearance refinement technique to suppress motion-induced artifacts, Catalyst4D enables high-quality and semantically meaningful edits across time. Extensive experiments demonstrate that it achieves superior visual fidelity compared to existing approaches.
Acknowledgements
@@
A2.3 Ablation on Anchor Construction
Anchors are formed from the intersection of cylinder-neighborhoods, rather than sampling individual points. An anchor is generated only when multiple nearby Gaussians consistently fall within an adaptively defined cylinder (Fig. A1). This design yields more reliable anchors for temporal propagation than naive sampling (e.g., FPS).
A2.4 Robustness of Sinkhorn Algorithm
A2.5 Comparison on the Monocular Dataset
@@
Figure A1: Comparison of anchor construction methods.
@@
$\min_{P\in\mathbb{R}_{+}^{n\times m}}\sum_{i=1}^{n}\sum_{j=1}^{m}D_{ij}P_{ij}-\lambda_{0}\sum_{i,j}P_{ij}\log P_{ij}\\ +\lambda_{1}\operatorname{KL}\!\left(\bm{P}\mathbf{1}_{m}\,\Big\|\,\frac{1}{n}\mathbf{1}_{n}\right)+\lambda_{2}\operatorname{KL}\!\left(\bm{P}^{\mathrm{T}}\mathbf{1}_{n}\,\Big\|\,\frac{1}{m}\mathbf{1}_{m}\right),$ (22)
s.t. $P_{ij}\geq 0,\forall i\in\{1,...,n\},j\in\{1,...,m\}$, where $\bm{P}$ is the transport plan matrix, $\lambda_{0}=0.1$, $\lambda_{1}=1.0$, $\lambda_{2}=1.0$ are regularization weights, and $\mathit{KL}$ denotes the Kullback-Leibler divergence.
From the optimal plan $\bm{P}^{*}$, we obtain the correspondence corrcorr by assigning each edited anchor to the source anchor with the highest transport mass:
$corr_{j}=\arg\max_{i\in\{1,\ldots,n\}}P^{*}_{ij}.$ (23)
@@
Notably, when computing the deformation of edited Gaussians, unlike the additive form used for positional deformation $\Delta\bm{\mu}$ and rotational deformation $\Delta\mathbf{q}$, scaling deformation $\Delta\mathbf{s}$ is defined multiplicatively to preserve proportional change:
${\bm{\mu}}^{t}$ $={\bm{\mu}}^{1}+\Delta{\bm{\mu}}^{t},$ (24)
@@
where $\Delta{\bm{\mu}}^{t}$, $\Delta\mathbf{q}^{t}$, $\Delta\mathbf{s}^{t}$ are the deformation quantities of edited Gaussians from frame 1 to frame $t$.
@@
For dynamic scene reconstruction with Swift4D, we adopt effective rank regularization [16] to suppress needle-like artifacts.
@@
Figure A5: Failure case (motion jitter illustration).

### 31 Mar 2026 v2

@@ 2 Related Work
The introduction of 3DGS [18, 26, 27, 28, 33] has addressed many of these limitations by enabling point-based scene representations with efficient rendering and flexible editing. Methods such as GaussiansEditors [7, 48] constrain gradient-based optimization to target 3D Gaussians, achieving efficient object-level editing while leveraging 2D diffusion models as priors. Subsequent works [6, 52, 49, 50] incorporate 3D structural information into diffusion-based editing to improve cross-view consistency, and others [1, 54, 10, 57] exploit depth information from 3DGS for geometry modification. High-quality stylization methods [29, 11, 60] extend image style transfer techniques to 3D scenes, further broadening 3DGS applications in artistic creation. Despite their capabilities, these approaches remain confined to static scenes. Building upon these strong 3D editing foundations, extending them to dynamic scenes with temporal coherence remains an open and compelling challenge.
@@ 3 Scene Representation
3D Gaussian Splatting. We adopt anisotropic 3D Gaussians [18] to represent the scene. Each Gaussian is defined by a center $\mu\in\mathbb{R}^{3}$ and a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, whose spatial influence at a point $\mathbf{x}$ is given by:
$G(\mathbf{x})=\exp\!\left(-\tfrac{1}{2}(\mathbf{x}-\mu)^{\mathrm{T}}{\Sigma}^{-1}(\mathbf{x}-{\mu})\right).$ (1)
The covariance is factorized as ${\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\mathrm{T}}\mathbf{R}^{\mathrm{T}}$, where $\mathbf{R}$ is a rotation matrix (parameterized by quaternion $\mathbf{q}$) and $\mathbf{S}$ is a scaling matrix (from scale vector $\mathbf{s}$). Each Gaussian additionally has an opacity $\sigma$ and spherical harmonics coefficients $\mathbf{sh}$ to encode view-dependent color. The full scene can thus be represented as a set: $\mathcal{G}=\{G_{i}:{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i},\sigma_{i},\mathbf{sh}_{i}\}_{i=1}^{N}$. During differentiable rendering, Gaussians are projected to 2D and composited via $\alpha$-blending. The pixel color $C$ along ray $\mathbf{r}$ is:
$C(\mathbf{r})=\sum_{i\in\mathcal{N}}\operatorname{SH}(\mathbf{sh}_{i},\mathbf{v})\alpha_{i}~\!\!\!\prod_{j=1}^{i-1}\!\!~(1-\alpha_{j}),$ (2)
@@
$\Delta{\mu},\Delta\mathbf{q},\Delta\mathbf{s}=\mathcal{F}_{\theta}({\mu}_{c},t;\mathbf{z}),$ (3)
where $\mathbf{z}$ is a scene-level latent code. The deformed parameters at time $t$ are computed as: ${\mu}_{t}={\mu}_{c}+\Delta{\mu}$, $\mathbf{q}_{t}=\mathbf{q}_{c}+\Delta\mathbf{q}$, and $\mathbf{s}_{t}=\mathbf{s}_{c}+\Delta\mathbf{s}$. This deformation-based formulation ensures temporal coherence and semantic consistency across frames. Our framework is compatible with various 4D Gaussian formulations, and in practice, we build upon Swift4D [53] and 4DGS [51], which serve as the base representations in our experiments.
@@ 4 Method
As illustrated in Fig. 2, Catalyst4D transfers 3D Gaussian edits from the canonical frame to the entire 4D sequence. Given edited first-frame Gaussians obtained by existing 3D Gaussian editors (e.g.[6, 19, 11]), our goal is to propagate the edits through time while maintaining geometric accuracy and temporal coherence. To this end, Catalyst4D introduces two key components: (1) Anchor-based Motion Guidance (AMG), which establishes region-level correspondences via anchors to enable localized, consistent deformation transfer; and (2) Color Uncertainty-guided Appearance Refinement (CUAR), which corrects occlusion-induced color artifacts by propagating confident appearance information along motion trajectories.
@@ 4.1 Anchor-based Motion Guidance
To establish robust anchor correspondences that account for varying densities and structural changes induced by editing, we compute a soft correspondence matrix $\mathbf{P}\in\mathbb{R}^{n\times m}$ using unbalanced optimal transport (UOT) [3, 44], solved via the Sinkhorn algorithm [8]. Here, each element $\mathbf{P}_{ij}$ represents the transport from edited anchor $\mathbf{p}_{j}^{\mathrm{edit}}$ to source anchor $\mathbf{p}_{i}^{\mathrm{src}}$, and reliable correspondences are written as:
$corr=\{(\mathbf{p}_{i}^{\mathrm{src}},\mathbf{p}_{j}^{\mathrm{edit}})\mid i=\operatorname*{arg\,max}_{k}\mathbf{P}_{kj}\}.$ (8)
@@
Specifically, at timestamp $t$, the position deformation $\Delta{\mu}_{\mathbf{g}}^{t}$ is computed as a weighted average of the corresponding source Gaussians’ deformations:
$\Delta{\mu}_{\mathbf{g}}^{t}=\frac{\sum_{\mathbf{g}^{\prime}\in\mathcal{G}^{\mathrm{1,sub}}_{\mathrm{src}}}w_{\mathbf{g}^{\prime}}\Delta{\mu}_{\mathbf{g}^{\prime}}^{t}}{\sum_{\mathbf{g}^{\prime}\in\mathcal{G}^{\mathrm{1,sub}}_{\mathrm{src}}}w_{\mathbf{g}^{\prime}}},$ (9)
where $\Delta{\mu}_{\mathbf{g}^{\prime}}^{t}$ is the temporal deformation of source Gaussian $\mathbf{g}^{\prime}$ from frame 1 to $t$, predicted by the deformation network. The weight $w_{\mathbf{g}^{\prime}}$ combines the Gaussian’s opacity $\sigma_{\mathbf{g}^{\prime}}$ and its Mahalanobis distance to $\mathbf{g}$:
$w_{\mathbf{g}^{\prime}}=\sigma_{\mathbf{g}^{\prime}}\exp\Big(-\frac{1}{2}({\mu}_{\mathbf{g}^{\prime}}-{\mu}_{\mathbf{g}})^{\mathrm{T}}{\Sigma}_{\mathbf{g}^{\prime}}^{-1}({\mu}_{\mathbf{g}^{\prime}}-{\mu}_{\mathbf{g}})\Big).$ (10)
@@ 4.2 Color Uncertainty-guided Appearance Refinement
Optical Flow Rendering. Given the temporal deformation $\Delta{\mu}^{t}$ from motion guidance, we first render optical flow maps describing the image-space displacement of each Gaussian from frame 1 to frame $t$. Inspired by [20, 62], for a viewpoint $v$, the 3D positions ${\mu}_{1}$ and ${\mu}_{t}$ are projected to 2D via the projection function $\operatorname{Proj}_{v}$, yielding the per-Gaussian displacement:
$f_{1\to t}^{v}=\operatorname{Proj}_{v}({\mu}_{t})-\operatorname{Proj}_{v}({\mu}_{1}).$ (11)
@@ 5.1 Experimental Setup
Dataset and Metrics. We conduct experiments on the multi-camera video datasets DyNeRF [25] and MeetRoom [23], as well as the monocular dataset HyperNeRF [36]. We compare our method with SOTA approaches: Instruct 4D-to-4D [35], CTRL-D [13], and Instruct-4DGS [22]. The results of all methods are evaluated on test views using CLIP [38] similarity and a temporal consistency score [15].
Implementation Details. We incorporate several baseline 3D editing methods, including the 3D-consistent editing approach DGE [6], the SDS inversion-based framework DreamCatalyst [19], and the high-resolution style transfer method SGSST [11]. For 4D scene representation, we employ Swift4D [53] under multi-camera settings, while utilizing 4DGS [51] for monocular scenes.
@@ 5.2 Qualitative Results
Qualitative comparison. The qualitative results achieved on different datasets are shown in Fig. 4, with the corresponding text prompts displayed on the leftmost side. As can be seen, Instruct 4D-to-4D [35] and Instruct-4DGS [22] lack fine-grained localization capabilities, often causing unintended modifications to non-target regions. Although CTRL-D [13] supports localized editing in dynamic scenes, its reliance on 2D image edits introduces limitations. When the 2D edits deviate significantly from the original scene content, the reconstructed 4D representation fails to accurately reflect them, resulting in degraded visual fidelity (see the 3rd and 5th rows of CTRL-D in Fig.4). In contrast, our 3D editing component employs iterative alignment of 2D and 3D representations, producing more natural and coherent edits. Building on these geometry-consistent 3D edits, Catalyst4D focuses on propagating them to 4D Gaussian scenes, achieving enhanced spatio-temporal consistency and visual realism.
@@ 5.3 Quantitative Results
Tab. 1 presents quantitative comparisons with Instruct 4D-to-4D (IN4D), Instruct-4DGS (I4DGS), and CTRL-D on two DyNeRF scenes (Sear-steak, Coffee-martini) and one MeetRoom scene (Trimming). Across all three scenes, our method achieves the highest CLIP similarity, indicating the strongest semantic alignment with the editing instructions. Regarding efficiency, our method outperforms CTRL-D and Instruct 4D-to-4D, and maintains a highly competitive training time comparable to Instruct-4DGS, while delivering vastly superior semantic fidelity. In terms of temporal coherence, measured by Consistency from VBench [15], our method also delivers consistently strong performance, matching or surpassing CTRL-D in most cases. We observe a slight drop in consistency on only one scene, which stems from our design choice of not retraining the deformation network or modifying the density of edited dynamic Gaussians. Despite this, the overall results demonstrate that our method maintains competitive stability while offering significantly better semantic fidelity.
@@
Figure 5: Comparison of anchor construction methods.
5.4 Ablation Studies
On Anchor Construction. An anchor is generated only when multiple nearby Gaussians consistently fall within an adaptively defined cylinder (Fig. 5). This design yields more reliable anchors for temporal propagation than naive sampling (e.g., FPS).
In contrast, AMG module establishes robust structural correspondences between edited and original Gaussians by matching sparse yet representative anchor points. This preserves regional motion and prevents interference from sparse noise, leading to improved geometric fidelity. The quantitative results in Tab. 2 confirm this observation, our method achieves higher CLIP similarity and better temporal consistency than relying on the deformation network (w/o AMG) to propagate the edited Gaussians.
@@ 6 Conclusion
We present Catalyst4D, a novel framework that brings the editing capabilities of 3D methods into the realm of dynamic 4D Gaussian scenes. By introducing an anchor-based motion guidance strategy to preserve structural consistency and a color uncertainty-guided appearance refinement technique to suppress motion-induced artifacts, Catalyst4D enables high-quality edits across time. Extensive experiments confirm its superior visual fidelity over existing approaches.
Acknowledgment
@@
A2.3 Comparison on Background Preservation
Although the compared CTRL-D [13] appears visually plausible, it still introduces undesired modifications in non-edited regions due to the limitations of the underlying 2D diffusion model. As illustrated in Fig.A1, while CTRL-D edits the character, it also causes noticeable deviations in non-target objects such as the dog on the stool and objects on the table compared to the original scene. By contrast, our method constrains target dynamic Gaussians directly through 3D editing gradients, enabling more precise and localized editing in dynamic scenes.
A2.4 Comparison on the Monocular Dataset
@@
Figure A1: Qualitative comparison of localized editing. In contrast to CTRL-D, which introduces inconsistencies in non-edited regions, our method achieves more precise and localized editing by constraining dynamic Gaussians via 3D editing gradients.
@@
$\min_{P\in\mathbb{R}_{+}^{n\times m}}\sum_{i=1}^{n}\sum_{j=1}^{m}D_{ij}\mathbf{P}_{ij}-\lambda_{0}\sum_{i,j}\mathbf{P}_{ij}\log\mathbf{P}_{ij}\\ +\lambda_{1}\operatorname{KL}\!\left(\mathbf{P}\mathbf{1}_{m}\,\Big\|\,\frac{1}{n}\mathbf{1}_{n}\right)+\lambda_{2}\operatorname{KL}\!\left(\mathbf{P}^{\mathrm{T}}\mathbf{1}_{n}\,\Big\|\,\frac{1}{m}\mathbf{1}_{m}\right),$ (22)
s.t. $\mathbf{P}_{ij}\geq 0,\forall i\in\{1,...,n\},j\in\{1,...,m\}$, where $\mathbf{P}$ is the transport plan matrix, $\lambda_{0}=0.1$, $\lambda_{1}=1.0$, $\lambda_{2}=1.0$ are regularization weights, and $\mathit{KL}$ denotes the Kullback-Leibler divergence.
From the optimal plan $\mathbf{P}^{*}$, we obtain the correspondence corrcorr by assigning each edited anchor to the source anchor with the highest transport mass:
$corr_{j}=\arg\max_{i\in\{1,\ldots,n\}}\mathbf{P}^{*}_{ij}.$ (23)
@@
Notably, when computing the deformation of edited Gaussians, unlike the additive form used for positional deformation $\Delta{\mu}$ and rotational deformation $\Delta\mathbf{q}$, scaling deformation $\Delta\mathbf{s}$ is defined multiplicatively to preserve proportional change:
${{\mu}}^{t}$ $={{\mu}}^{1}+\Delta{{\mu}}^{t},$ (24)
@@
where $\Delta{{\mu}}^{t}$, $\Delta\mathbf{q}^{t}$, $\Delta\mathbf{s}^{t}$ are the deformation quantities of edited Gaussians from frame 1 to frame $t$.
@@
For dynamic scene reconstruction with Swift4D, we adopt effective rank regularization [16] to suppress needle-like artifacts. Following the settings of 4DGS [51], we constrain the edited dynamic Gaussians to share consistent opacity and color attributes across all frames during optimization.
@@
Figure A5: Failure case (motion jitter illustration).